### v2

# Training-Free Text-Guided Color Editing with Multi-Modal Diffusion Transformer

Abstract.
Text-guided color editing in images and videos is a fundamental yet unsolved problem, requiring fine-grained manipulation of color attributes, including albedo, light source color, and ambient lighting, while preserving physical consistency in geometry, material properties, and light-matter interactions. Existing training-free methods offer broad applicability across editing tasks but struggle with precise color control and often introduce visual inconsistency in both edited and non-edited regions. In this work, we present ColorCtrl, a training-free color editing method that leverages the attention mechanisms of modern Multi-Modal Diffusion Transformers (MM-DiT). By disentangling structure and color through targeted manipulation of attention maps and value tokens, our method enables accurate and consistent color editing, along with word-level control of attribute intensity. Our method modifies only the intended regions specified by the prompt, leaving unrelated areas untouched. Extensive experiments on both SD3 and FLUX.1-dev demonstrate that ColorCtrl outperforms existing training-free approaches and achieves state-of-the-art performances in both edit quality and consistency. Furthermore, our method surpasses strong commercial models such as FLUX.1 Kontext Max and GPT-4o Image Generation in terms of consistency. When extended to video models like CogVideoX, our approach exhibits greater advantages, particularly in maintaining temporal coherence and editing stability. Finally, our method also generalizes to instruction-based editing diffusion models such as Step1X-Edit and FLUX.1 Kontext dev, further demonstrating its versatility. Here is the project website.
copyright: nonecopyright: none
Figure 1. Text-conditioned color editing. Our method, ColorCtrl with FLUX.1-dev, edits colors across multiple materials while preserving light-matter interactions. For example, in the fourth case, the ball’s color, its water reflection, specular highlights, and even small droplets on the glass have all been changed. It also enables fine-grained control over the intensity of specific descriptive terms.
1. Introduction
Film industry-level color changing in images and videos based on textual instructions is a fundamental yet challenging task in deep learning and visual editing. Here, “color” encompasses not only object albedo (i.e., intrinsic surface color independent of material properties), but also the color of light sources and ambient illumination. Color editing is an ill-conditioned problem, as it requires explicit or implicit 3D reconstruction of the entire scene, including correct illumination. During editing, it is essential to modify only the intended color attributes while preserving material properties, ensuring accurate reflections and refractions, and keeping non-editing regions unchanged.
Traditional image processing methods have been widely commercialized in professional software such as Photoshop, serving billions of users worldwide. However, the steep learning curve and significant manual effort required make it difficult to achieve widespread accessibility. Moreover, such tools are not well-suited for automated batch processing and present challenges when applied to video-related tasks. Recently, diffusion models have demonstrated remarkable capabilities in generating high-quality images that adhere to physical principles of color and illumination. This has spurred growing interest in leveraging their generative power to tackle the above challenges, with controllability emerging as a critical factor. Although many methods (Magar et al., 2025; Zhang et al., 2025) fine-tune diffusion models for controllable editing, they usually require large-scale datasets and complex training pipelines, and are often constrained to narrow domains or specific edit types. On the other hand, training-free methods (Jiao et al., 2025) have gained popularity for a wide range of image editing tasks due to their generality and ease of use. Despite their success in many scenarios, they still struggle with fine-grained color control and often introduce inconsistencies in edited and non-edited regions, making precise color editing an unresolved challenge.
The recent architectural shift in diffusion models from U-Net (Rombach et al., 2022) to the Multi-Modal Diffusion Transformer (MM-DiT) (Esser et al., 2024) opens up new opportunities for training-free color editing, for two main reasons: (1) the new architecture allows for scaling up both data and model capacity, enabling better adherence to physical priors; and (2) the improved fusion of text and vision modalities supports more flexible and precise attention-control editing strategies.
In this work, we propose ColorCtrl, a training-free, open-world color editing method that effectively leverages pre-trained MM-DiT models to perform natural and precise color modifications, while preserving all other visual attributes, such as geometry, material, reflection and refraction, light source position, illumination direction, and light intensity, as shown in Fig. 1. For example, when editing the color of a ball, ColorCtrl accurately adjusts not only the ball itself but also its reflection in the water, the specular reflections on both sides, and even the small water droplets on the glass surface, as demonstrated in the fourth example of Fig. 1. Furthermore, ColorCtrl supports fine-grained, word-level control over the strength of color attributes while preserving the other visual attributes mentioned above. Thanks to our precise and robust control, ColorCtrl requires no manual parameter tuning and can be directly applied across all attention layers and inference timesteps.
To verify the effectiveness of our method, we conduct extensive experiments showing that ColorCtrl outperforms existing training-free approaches on both Stable Diffusion 3 Medium (a.k.a., SD3) (Esser et al., 2024) and FLUX.1-dev (Labs, 2024). To better understand the gap between open-source and commercial models, we further compare ColorCtrl with two strong commercial baselines: FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (a.k.a., GPT-4o) (OpenAI, 2025). ColorCtrl achieves more natural color editing and substantially better preservation of visual consistency, including background and structural fidelity. In addition, our method is model-agnostic and can be seamlessly extended to video models such as CogVideoX (Yang et al., 2024), where the performance gap becomes even more pronounced. Finally, ColorCtrl can be integrated into instruction-based editing models such as Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025), demonstrating strong compatibility.
In summary, the main contributions can be listed as follows.
  • We propose ColorCtrl, a training-free method for color editing that supports modification of albedo, light source color and ambient lighting, while preserving the consistency of geometry, material and light-matter interaction.
  • We present extensive experiments demonstrating that ColorCtrl achieves state-of-the-art performance among training-free methods on MM-DiT-based models. Compared to commercial models (i.e., FLUX.1 Kontext Max and GPT-4o Image Generation), our method delivers significantly better consistency preservation and produces more natural edits.
  • ColorCtrl generalizes well across multiple MM-DiT-based models, including video and instruction-based editing models, highlighting its broad applicability and extensibility.
2. Related Work
Text-to-image and video generation. Diffusion models with a U-Net backbone (Ho et al., 2020; Rombach et al., 2022; Guo et al., 2024) have largely replaced early GAN systems (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) due to superior image fidelity. However, the U-Net design scales poorly, leading to the adoption of Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among these, MM-DiT (Esser et al., 2024) has emerged as a widely adopted backbone for recent state-of-the-art models (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025b; Black Forest Labs, 2025), such as SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, as well as CogVideoX (Yang et al., 2024) for video generation. In this work, we propose an attention-control scheme that plugs into any MM-DiT model.
Text-guided editing. Training-free text-guided editing methods based on pre-trained diffusion models offer flexibility and efficiency. Existing approaches fall into two groups: (i) sampling-based methods, which steer generation by injecting controlled noise or inversion (Jiao et al., 2025; Huberman-Spiegelglas et al., 2024; Xu et al., 2023; Kulikov et al., 2024; Yan et al., 2025); (ii) attention-based methods, which modify attention maps, starting with Prompt-to-Prompt (Hertz et al., 2023) and its image/video variants (Chen et al., 2024; Wang et al., 2025; Liu et al., 2024; Cao et al., 2023; Cai et al., 2025; Rout et al., 2025; Xu et al., 2025). Among these, DiTCtrl is the only method exploring attention control in MM-DiT, but ColorCtrl differs in key ways: DiTCtrl only applies mask extraction during long video generation, not color editing, while our improved method is more robust. Additionally, DiTCtrl’s attribute re-weighting disrupts attention consistency, causing inconsistent geometries, while our method avoids this. Furthermore, ColorCtrl works across all layers and timesteps, unlike DiTCtrl, which requires careful layer selection. Different from DiTCtrl, we operate in attention maps rather than key and value tokens. More recently, methods (Liu et al., 2025a; Brooks et al., 2023) like FLUX.1 Kontext (Black Forest Labs, 2025) and GPT-4o (OpenAI, 2025) have shown convenient creative workflows in training with synthetic instruction-response pairs for fine-tuning a diffusion model for image editing.
Despite advancements in text-guided editing, we demonstrate that current methods still face significant challenges in accurately changing colors while preserving geometry, material properties and light-matter interaction. Also, all aforementioned training-free methods rely on selectively manipulating specific inference steps or attention layers, which limits their robustness and consistency with respect to the source. In contrast, our approach requires no manual selection of steps or layers.
3. Method
We first formulate the color editing problem in Sec. 3.1, followed by a revisit of the attention mechanism in MM-DiT blocks in Sec. 3.2. Sec. 3.3 describes our approach for preserving geometry, material properties, and light-matter interactions. In Sec. 3.4, we introduce a method for preserving colors in non-editing regions, which are automatically detected by the model. Together, Sec. 3.3 and Sec. 3.4 constitute the core of our method. Finally, Sec. 3.5 introduces word-level control over color attributes via attention re-weighting, serving as an optional add-on for fine-grained attribute control.
3.1. Task Formulation: Text-guided Color Editing
Rendering process. To provide a clearer and more precise description of the conditions required for film industry-level color changing, we define the rendering process of a frame by
(1) $\textbf{I}\;=\;\mathcal{R}\bigl{(}G,\;L,\;A,\;S,\;C\bigr{)},$
where the notations of rendering inputs are defined as,
  • $\mathcal{R}(\cdot)$ - the rendering function,
  • $G=\{G_{1},\dots,G_{N}\}$ - a set of object geometries, e.g., mesh topology for $N$ objects,
  • $L=\{L_{\text{env}},L_{1},\dots,L_{K}\}$ - the ambient illumination $L_{\text{env}}$ and $K$ light sources (each with position, intensity, and spectrum),
  • $A=\{A_{1},\dots,A_{N}\}$ - the albedos (base colors) of each object,
  • $S=\{S_{1},\dots,S_{N}\}$ - material parameters (roughness, specularity, and normal maps) for each object that are color-independent,
  • $C$ - camera intrinsics and extrinsics.
Editing task. According to the predefined image rendering process in Eq. 1, we formulate the consistent editing task in this work as follows. Given a source image I and text prompt pairs $q$ before and after editing [^1] An example of $q$: “white fox.” $\to$ “orange fox.” (source and target prompt respectively). that specifies which objects or lights to recolor and the desired target color, we aim to learn the following editing function $f(\cdot)$
(2) $f:(\textbf{I},q)\,\longmapsto\,\hat{\textbf{I}}=\mathcal{R}\bigl{(}G,\hat{L},\hat{A},S,C\bigr{)},$
such that $\hat{\textbf{I}}$ satisfies:
  1. (C1)
    Geometry/view consistency - $G$ and $C$ remain fixed with the source image I, preserving object layout and perspective.
  2. (C2)
    Illumination consistency - light positions and scalar intensities remain fixed. Specifically, changes may occur only on the target spectral components (i.e., RGB channels) of $L$.
  3. (C3)
    Material consistency - the object material $S$ remain fixed. The edits apply only to the albedo components $A$ of specific objects.
In the scope of the defined editing task in Eq. 2, the editing function $f(\cdot)$ needs to internally infer the underlying scene parameters and localize the editing objects, lights, and regions to perform precise, constrained color changes.
Figure 2. Pipeline of ColorCtrl. (a) Visualizes the attention mechanism in MM-DiT blocks. (b) Enables effective color editing while maintaining structural consistency. (c) Preserves colors in non-editing regions. (d) Applies attribute re-weighting to specific word tokens. Symbols in the source branch have no superscript. Symbols with a superscript indicate the target branch, and hats (e.g., $\hat{V}$, $\hat{M}$) denote edited outputs.
3.2. Preliminaries: Attention in MM-DiT Blocks
The attention mechanism for text-vision fusion in MM-DiT (Esser et al., 2024) differs from that in U-Net (Rombach et al., 2022). In U-Net, cross-attention is used for text guidance, while self-attention focuses on visual content interactions, separately. In contrast, MM-DiT integrates text and vision by concatenating their tokens together and processing them jointly via sole self-attention. As illustrated in Fig. 2 (a), vision and text tokens $\bm{z}$ are fed into the $i$-th MM-DiT block at timestep $t$ during inference. After modulation, the block produces an attention map $\bm{M}$ and value tokens $\bm{V}$, which are used to generate the updated tokens $\hat{z}$ for the next block and timestep. The resulting attention map $\bm{M}$ can be divided into four parts: vision-to-vision (upper-left), vision-to-text (upper-right), text-to-vision (lower-left), and text-to-text (lower-right). These quadrants correspond to different query-key token pairings. For example, the text-to-vision region represents attention scores computed between query tokens from the text modality $\bm{Q^{\text{text}}}$ and key tokens from the vision modality $\bm{K^{\text{vision}}}$. Similarly, the value tokens $\bm{V}$ are composed of a vision part $\bm{V^{\text{vision}}}$ and a text part $\bm{V^{\text{text}}}$. The functional roles of each region are discussed in detail in the following sections.
3.3. Structure Preservation
Following (Hertz et al., 2023; Cao et al., 2023), we divide the editing process into a source branch and a target branch. The source branch follows the original generation process, producing a source image and storing intermediate attention outputs. The target branch reuses these stored variables to generate edited results. Starting from a fixed random seed, the model can generate the desired object without applying any editing method. However, the resulting image often diverges significantly from the source, making it unsuitable for meaningful editing tasks that require structural consistency. To address this, all subsequent modifications are applied exclusively on the target branch, aiming to improve consistency without disrupting the intended edit. The process of keeping $G$, $S$, $C$, light positions, and scalar intensities fixed (as in constraints (C1)-(C3)) is referred to as structure preservation. We observe that the vision-to-vision part of the attention map inherently encodes rich knowledge about the parts of the scene that must remain unchanged. Given a source attention map $\bm{M}$, its vision-to-vision part is transferred from $\bm{M}$ to the corresponding part in the target attention map $\bm{M^{*}}$, producing an edited attention map $\bm{\hat{M}}$ that fully respects the structure preservation constraints, as described in Fig. 2 (b).
3.4. Color Preservation
   Method    SD3    FLUX.1-dev
   Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$    Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$
   $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited    $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited
   Fix seed    0.5787    20.44    0.8411    29.17    27.54    0.7180    22.32    0.8877    27.72    25.76
   FireFlow (Deng et al., 2025)    0.6078    19.19    0.8461    28.63    27.24    0.8322    35.87    0.9717    25.84    23.56
   RF-Solver (Wang et al., 2025)    0.6711    23.30    0.8906    27.43    25.96    0.8394    36.25    0.9715    25.83    23.58
   SDEdit (Meng et al., 2022)    0.6353    27.78    0.8699    25.98    24.47    0.8285    32.62    0.9605    25.54    23.17
   DiTCtrl (Cai et al., 2025)    0.8119    35.40    0.9812    26.21    24.67    0.8306    34.48    0.9791    25.89    23.58
   UniEdit-Flow (Jiao et al., 2025)    0.8016    36.31    0.9774    26.08    24.67    0.8498    37.57    0.9777    25.78    23.44
   Ours    0.8473    42.93    0.9960    28.32    26.96    0.9196    39.49    0.9936    27.34    24.90
Table 1. Quantitative image results compared with training-free methods on PIE-Bench. Results for FireFlow on SD3 are omitted due to their consistency being worse than that obtained using fixed seeds.
Figure 3. Top row: SD3 results; bottom row: FLUX.1-dev results. (a) The edit prompt is “white fox” $\to$ “orange fox”. Left to right: source image, our full method, without color preservation, with swapped text-to-text part in structure preservation, and with swapped $V^{\text{text}}$ in color preservation. (b) The generation prompt is “a white fox in a forest”, and the token for mask extraction is “fox”. From left to right: the mask extracted from vision-to-text parts, and from text-to-vision parts.
Despite enforcing structure preservation, we observe that undesired changes such as color shifts can still occur in regions unrelated to the edit, as shown in Fig. 3 (a). To further localize the edit and reduce inconsistencies, edits should be confined to the intended regions, while preserving all other areas. Motivated by the approach of Cai et al. (2025), we first extract a binary mask $\bm{m}$ from vision-to-text parts of attention maps with a threshold $\epsilon$, which indicates the target editing region. In contrast to Cai et al. (2025), which averages the vision-to-text and text-to-vision parts to obtain the mask, we use only the vision-to-text parts, as they provide superior spatial localization for the target object, unlike the text-to-vision parts (see Fig. 3 (b)). Based on $\bm{m}$, we copy the value tokens from the non-editing regions in the vision part of the source $\bm{V}$ into the corresponding regions of the target $\bm{V^{*}}$, yielding the final value tokens $\bm{\hat{V}}$, as shwon in Fig. 2 (c). We refer to this procedure as color preservation. As shown in Fig. 3 (a), including the text part of value tokens $\bm{V^{\text{text}}}$ during value exchange significantly weakens text guidance in FLUX.1-dev, and leads to severe artifacts in SD3.
3.5. Attribute Re-Weighting
So far, our method enables powerful color editing capabilities. However, the granularity of control via text prompts remains limited. For example, if the user specifies a color like “dark yellow”, there is no way to explicitly control the degree of darkness. Existing re-weighting techniques in U-Net-based models primarily follow two approaches: (1) Scaling the text embedding of specific tokens before the diffusion process (Perry, 2023). However, this method is designed for CLIP-based (Radford et al., 2021) text encoders and is incompatible with MM-DiT-based models, which typically use T5 (Raffel et al., 2020). (2) Scaling the attention scores of specific tokens in the cross-attention map (Hertz et al., 2023). This is also inapplicable to MM-DiT, which relies solely on self-attention. Moreover, both approaches fail to maintain structural consistency during scaling, limiting their utility in high-fidelity editing tasks.
To support more fine-grained and user-friendly control in MM-DiT-based models, we introduce a mechanism to modulate the strength of specific attribute words. Specifically, we scale the attention scores in the text-to-vision parts of the attention map corresponding to the selected word tokens before the softmax operation, as illustrated in Fig. 2 (d). This modification can be applied either directly on the original source attention maps or on the target attention maps that have already undergone structure preservation, allowing flexible integration in above editing pipeline.
4. Experiments
4.1. Setup
Baselines. We compare our method against several training-free approaches built upon MM-DiT, including FireFlow (Deng et al., 2025), RF-Solver (Wang et al., 2025), SDEdit (Meng et al., 2022), DiTCtrl (Cai et al., 2025), FlowEdit (Kulikov et al., 2024), and UniEdit-Flow (Jiao et al., 2025). We focus exclusively on MM-DiT-based baselines, as prior work (Jiao et al., 2025; Deng et al., 2025) has shown that U-Net-based methods perform significantly worse. Accordingly, we exclude methods that cannot be adapted to the MM-DiT architecture. For commercial baselines, we compare with FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025).
Implementation. We conduct experiments on SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, and on CogVideoX-2B (Yang et al., 2024) for video generation. For instruction-based image editing, we use Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025). For FLUX.1-dev, Step1X-Edit, and FLUX.1 Kontext dev, we apply attention control to the single-stream attention layers, following Deng et al. (2025). Unless otherwise noted, we use the Euler sampler and adopt UniEdit-Flow (Jiao et al., 2025) for image inversion. Maintaining a balance between fidelity to the original image and the strength of the applied edits is a well-known trade-off in generative editing. To ensure a fair comparison across methods, we carefully tune the hyperparameters for each baseline. Additional details are provided in Appendix A.
Figure 4. Qualitative image results compared with training-free methods and commercial models on PIE-Bench. The top three rows are generated using FLUX.1-dev, while the bottom two are generated using SD3. Best viewed with zoom-in.
   Method    SD3    FLUX.1-dev
   Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$    Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$
   $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited    $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited
   FLUX.1 Kontext Max (Black Forest Labs, 2025)    0.7305    31.97    0.9254    28.93    27.30    0.7607    26.77    0.9165    28.36    26.10
   GPT-4o Image Generation (OpenAI, 2025)    0.6240    23.69    0.8134    29.51    28.08    0.7431    23.71    0.8755    28.84    26.46
   Ours    0.8473    42.93    0.9960    28.32    26.96    0.9196    39.49    0.9936    27.34    24.90
Table 2. Quantitative image results compared with commercial models on PIE-Bench.
Benchmark. While prior editing methods (Hertz et al., 2023; Cao et al., 2023; Cai et al., 2025) typically lack standardized benchmark evaluation, we adopt prompts from the Change Color task in PIE-Bench (Ju et al., 2024), which includes 40 editing pairs, to better showcase the capabilities of our method. Although our approach is fully compatible with inversion methods, we adopt a noise-to-image setting to better isolate and evaluate editing performance, removing the influence of reconstruction and inversion. This also allows us to reuse the same set of benchmark prompts for video diffusion models, enabling a consistent and fair comparison across both image and video domains. To further scale up the evaluation, we introduce a new benchmark called ColorCtrl-Bench, consisting of 300 prompt pairs in the same style as PIE-Bench (see Sec. A.4 for details). For all baselines, we adopt a fixed sampler and identical random seeds to ensure that source images are consistent, enabling reliable comparison across methods.
Evaluation protocol. Unlike the original PIE-Bench, which evaluates structural similarity using structural distance (Tumanyan et al., 2022), we adopt the Structural Similarity Index (SSIM) (Wang et al., 2004) computed on Canny edge maps (Canny, 1986), following the approach of Zhao et al. (2023), for more accurate assessment. To evaluate the preservation of non-edited regions (a.k.a., BG Preservation), we compute PSNR and SSIM exclusively on those regions, which are annotated using Grounded SAM 2 (Ren et al., 2024) with dilation. Semantic alignment of the edits is assessed using CLIP similarity (Radford et al., 2021), applied to both the entire image and the edited regions.
4.2. Comparison with Training-Free Methods (Images)
Tab. 1 and Tab. 5 report benchmark results on both SD3 and FLUX.1-dev, comparing our method with other training-free baselines. Our method achieves state-of-the-art performance, delivering superior results in both preserving source content and executing accurate edits. Fig. 4 further supports this finding: other methods exhibit limited capacity for color editing and often introduce visual inconsistencies, while ours produces coherent and faithful edits. Additional results are in Appendix B.
4.3. Comparison with Commercial Models (Images)
Tab. 2 and Tab. 6 compare our method (based on SD3 and FLUX.1-dev) with two commercial models: FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025). Despite being based on open-source models, our approach achieves superior layout and detail consistency, as well as better preservation of non-edited regions. While the CLIP similarity scores of our method are slightly lower, visual results in Fig. 4 reveal that the commercial models often rely on over-saturated, unrealistic edits to better align with prompts. For example, in the top row, FLUX.1 Kontext Max recolors the entire mouse, including its magic wand, in solid purple, while GPT-4o produces a dark, dissonant shade. In contrast, our method applies a harmonious color. In the second row, only our method preserves the shirt’s semi-transparency, whereas the commercial models render it as opaque yellow, ignoring material properties. In the third row, our method respects the muted tone of “green tea” and edits the ice cream to a natural reddish-brown “red tea” color. The commercial models, however, apply an unnaturally pure red. In the final row, although the prompt requests a “yellow kitten”, no naturally occurring cat has a truly pure yellow coat. Our method instead generates a kitten with the closest plausible fur color, aligned with real-world appearances, unlike the commercial models, which apply flat, high-saturation tones that appear unnatural. These results demonstrate that higher CLIP similarity does not necessarily indicate better edit quality. It often stems from prompt overfitting while compromising realism and consistency. Overall, our method consistently produces more faithful, controllable edits, even built on open-source models.
4.4. Comparison with Training-Free Methods (Videos)
   Method    PIE-Bench (Ju et al., 2024)    ColorCtrl-Bench
   Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$    Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$
   $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited    $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited
   Fix seed    0.6228    20.73    0.8912    27.97    26.63    0.6671    19.49    0.8489    28.01    26.66
   FireFlow (Deng et al., 2025)    0.7517    30.74    0.9605    25.42    24.18    0.7957    28.78    0.9323    25.26    24.12
   RF-Solver (Wang et al., 2025)    0.7677    35.41    0.9730    25.40    24.08    0.8102    32.64    0.9562    25.36    24.16
   SDEdit (Meng et al., 2022)    0.7880    35.46    0.9689    24.64    23.21    0.8281    34.30    0.9626    24.74    23.46
   DiTCtrl (Cai et al., 2025)    0.6912    27.70    0.9435    27.02    25.57    0.7270    26.14    0.9165    26.54    25.14
   FlowEdit (Kulikov et al., 2024)    0.7852    34.24    0.9704    26.13    24.97    0.8639    32.64    0.9774    25.95    23.63
   UniEdit-Flow (Jiao et al., 2025)    0.7645    38.00    0.9772    25.35    24.13    0.8002    35.33    0.9613    25.39    24.14
   Ours    0.8651    38.98    0.9916    27.12    25.96    0.8885    38.68    0.9893    27.32    26.16
Table 3. Quantitative video results compared with baselines on PIE-Bench and ColorCtrl-Bench.
Figure 5. Qualitative video results compared with training-free methods on PIE-Bench. Prompt is “green tea” $\to$ “red tea”. Each shows three frames.
Thanks to the training-free nature of our method, it can be seamlessly extended to video editing tasks. Since FLUX.1 Kontext Max and GPT-4o do not support video editing, we compare only with other training-free methods. FlowEdit is excluded as it only applies to rectified flow models rather than diffusion-based models. Tab. 3 presents benchmark results on CogVideoX-2B (Yang et al., 2024). Similar to image editing, our method outperforms all baselines. Notably, the performance gap becomes even more pronounced due to the added temporal dimension. Visualization results in Fig. 5 further highlight the effectiveness of ColorCtrl. Consistent with its performance in image editing, our method also handles challenging cases in video, such as accurately reflecting the color change of the ice cream in the bowl’s reflection. Please refer to Appendix B for additional comparisons.
4.5. Attribute Re-Weighting Analysis
Figure 6. Examples of attribute re-weighting. The top two rows are generated using FLUX.1-dev, while the bottom one are generated using SD3.
Although DiTCtrl (Cai et al., 2025) includes a mechanism for attribute re-weighting, both the paper and results in Tab. 1 suggest limitations. Specifically, the method struggles to achieve the intended color changes while maintaining consistency with the source image. Moreover, their approach of scaling attention weights after the softmax operation violates the assumption that attention scores should sum to one, which leads to incorrect attention behavior. Since none of the baselines can smoothly adjust attribute strength while simultaneously satisfying the three constraints ((C1)-(C3)), we present results of our method alone in Fig. 6. These results show that our method not only supports re-weighting a single attribute within the same prompt, but also allows adjusting attribute strength across different prompts (second row). Moreover, our method can re-weight multiple attributes simultaneously (third row). Overall, these results demonstrate that our method enables smooth and controllable transitions in attribute strength, while preserving structural consistency across the image and maintaining color fidelity in non-edited regions, on both SD3 and FLUX.1-dev.
4.6. Ablation Study
According to Tab. 4, starting from fixed random seed generation, we observe the highest CLIP similarity due to the lack of consistency constraints. However, this comes at the cost of very low scores in Canny SSIM, as well as PSNR and SSIM in non-edited regions, indicating poor structural and visual consistency. Introducing the structure preservation component significantly improves all consistency metrics, suggesting that the geometric and material attributes are effectively maintained, as also illustrated in Fig. 3. Adding the color preservation component completes our method, which further enhances consistency, particularly in non-edited regions, while sacrificing almost no CLIP similarity. Overall, the results validate the effectiveness of each component in our method.
Figure 7. Examples of results generated with Step1X-Edit (left) and FLUX.1 Kontext dev (right). Green arrows: first edit using the editing model. Blue arrows: second edit using the editing model. Orange arrows: second edit using the editing model with ColorCtrl. Top left: a red diamond is added to the neck, then changed to blue. Bottom left: an orange cap is added, then changed to purple. Top right: blue butterflies are added, then changed to yellow. Bottom right: a flashlight with blue light is added, then changed to green.
   Method    SD3    FLUX.1-dev
   Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$    Canny    BG Preservation    $\text{Clip Similarity}$$\uparrow$
   $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited    $\text{SSIM}$$\uparrow$    PSNR $\uparrow$    SSIM $\uparrow$    Whole    Edited
   Fix seed    0.5787    20.44    0.8411    29.17    27.54    0.7180    22.32    0.8877    27.72    25.76
    + Structure Preservation    0.7312    24.77    0.9201    28.41    27.29    0.9019    30.20    0.9719    27.44    25.22
    + Color Preservation (Ours)    0.8473    42.93    0.9960    28.32    26.96    0.9196    39.49    0.9936    27.34    24.90
Table 4. Ablation study evaluating the effectiveness of each component on PIE-Bench.
4.7. Compatibility and Discussion
Real image editing.
Figure 8. Examples of real image editing. Left: results generated with SD3. Right: results generated with FLUX.1-dev.
To apply ColorCtrl to real images, we integrate UniEdit-Flow (Jiao et al., 2025) for inversion and replace the original editing module with our method. The results, shown in Fig. 8, demonstrate that our approach generalizes well to real-world inputs on both SD3 and FLUX.1-dev, preserving fine-grained consistency (e.g., subtle fabric wrinkles and shadows) while delivering strong editing performance. Notably, in the top row, even when editing black clothing, ColorCtrl accurately distinguishes material shading from cast shadows, resulting in illumination-consistent edits.
Generalization to instruction-based editing diffusion models. In addition to text-to-image and text-to-video models, ColorCtrl is also compatible with instruction-based editing diffusion models, such as Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025). Given a real input image and a target editing instruction, the model performs edits accordingly. However, performing a second round of editing using the original model alone often leads to structural inconsistencies, such as distortions or shifts in shadows and edges. By incorporating our method, the model can further refine color edits while preserving structural fidelity. As shown in Fig. 7, compared to using the base model alone, our approach achieves better outline consistency and improved preservation of subtle visual cues like shadows.
Discussion. Our method integrates structure preservation, regional color preservation, and word-level attribute intensity control into a unified, training-free pipeline. These components work together to enable precise and consistent text-driven color editing. Each part of the attention computation plays a distinct role: the vision-to-vision part of $\bm{M}$ preserves structure, the vision-to-text part is used for mask extraction, the text-to-vision part enables attribute re-weighting, the vision part of $\bm{V}$ supports color preservation, and the text-to-text region of $\bm{M}$ along with the text part of $\bm{V}$ provides crucial textual guidance. As demonstrated by the degradation observed in Fig. 3 (a) when the text-to-text region of $\bm{M}$ or the text part of $\bm{V}$ is altered, preserving the integrity of these parts is essential for maintaining robust and coherent generation that aligns with the target prompt.
5. Conclusion
We have introduced ColorCtrl, a training-free method for text-guided color editing that achieves fine-grained, physically consistent control over albedo, light source color, and ambient illumination. Our method is designed to edit only the intended visual attributes specified in the prompt, leaving all unrelated regions untouched. Built upon diverse MM-DiT-based diffusion models, such as SD3 and FLUX.1-dev, our approach enables fine-grained control over attribute intensity, while preserving geometry, material properties, and light-matter interactions. ColorCtrl not only outperforms prior training-free methods and achieves state-of-the-art results but also delivers stronger consistency than commercial models, i.e., FLUX.1 Kontext Max and GPT-4o Image Generation, on both quantitative and qualitative evaluations. Moreover, its model-agnostic design generalizes naturally to video diffusion models (i.e., CogVideoX) and instruction-based editing diffusion models (i.e., Step1X-Edit and FLUX.1 Kontext dev), highlighting its broad applicability. We believe ColorCtrl paves the way for scalable, high-fidelity, and controllable color editing in both research and practical deployment.
Appendix A Implementation Details
A.1. Inference Settings
For benchmarking, we use 28 inference steps for both SD3 (Esser et al., 2024) and FLUX.1-dev, with the classifier-free guidance (CFG) scale (Ho and Salimans, 2021) set to 7.5. All images are generated at a resolution of 1024 $\times$ 1024. For CogVideoX-2B, we use 50 inference steps and set the CFG scale to 6, generating videos at 49 frames with 720 $\times$ 480 resolution. A fixed random seed of 42 is used for all benchmark experiments. For real image editing, we use the latest inversion method from UniEdit-Flow (Jiao et al., 2025) and set the CFG scale to 1 according to the method.
The target object for mask extraction is determined using the “blended_word” keywords provided by PIE-Bench (Ju et al., 2024) and ColorCtrl. We use a mask threshold of $\epsilon=0.1$, which consistently yields strong performance across SD3, FLUX.1-dev, and CogVideoX-2B. Despite being relatively coarse, this threshold is sufficient due to the strong global adaptation ability of modern generative models, which enables them to propagate edits from sparse color cues to semantically aligned regions.
For Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025), we use their official code and setting with resolution 1024 $\times$ 1024.
Image generation is performed on an RTX 4090 GPU, while video generation uses an A100 GPU.
A.2. Sampling Details
Method SD3 FLUX.1-dev
Canny BG Preservation $\text{Clip Similarity}$$\uparrow$ Canny BG Preservation $\text{Clip Similarity}$$\uparrow$
$\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited $\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited
Fix seed 0.6805 16.93 0.8303 28.27 26.65 0.7569 19.54 0.8552 27.70 26.25
FireFlow (Deng et al., 2025) 0.6625 13.87 0.7664 27.45 25.95 0.8553 32.80 0.9594 25.05 23.69
RF-Solver (Wang et al., 2025) 0.7329 18.69 0.8596 26.70 25.19 0.8832 36.25 0.9736 24.89 23.49
SDEdit (Meng et al., 2022) 0.7643 27.24 0.9287 25.22 23.77 0.8430 29.47 0.9428 24.45 22.83
DiTCtrl (Cai et al., 2025) 0.8465 31.65 0.9723 25.25 23.82 0.8699 31.72 0.9707 24.90 23.53
FlowEdit (Kulikov et al., 2024) 0.8261 30.77 0.9622 25.53 24.24 0.8704 29.29 0.9620 25.09 23.52
UniEdit-Flow (Jiao et al., 2025) 0.8503 32.74 0.9725 25.24 23.81 0.8725 34.85 0.9676 25.04 23.61
Ours 0.8775 38.16 0.9896 28.07 26.69 0.9324 37.96 0.9901 26.53 25.26
Table 5. Quantitative image results compared with training-free methods on ColorCtrl-Bench. Results for FireFlow on SD3 are omitted due to consistency worse than that obtained using fixed seeds.
To avoid redundant computation and accelerate inference, the source branch is first executed to cache attention maps and value tokens for reuse, following a similar strategy to that in Wang et al. (2025). During this stage, the editing mask $m$ is also computed to indicate the regions to be modified. In the actual editing phase, the cached features and mask are directly loaded, eliminating the need to recompute the source branch. This approach ensures that the editing process remains as efficient as standard sampling methods, without introducing any additional computational cost.
A.3. Implementation of Compared Methods
Several compared methods lack official implementations for SD3 or CogVideoX-2B, or do not provide compatible sampling code. To ensure fair comparison, we reimplement these methods on SD3, FLUX.1-dev, and CogVideoX-2B by faithfully following their original designs and carefully tuning hyperparameters to match the reported performance. Detailed implementations are as follows:
Figure 9. Examples of FireFlow and RF-Solver with difference end timesteps on SD3. The selected end timestep refers to the setting used in the benchmark evaluation, while the higher end timestep denotes a larger value chosen for comparison purposes.
  • DiTCtrl (Cai et al., 2025): For SD3-based image editing, we set the editing range from timestep $2$ to $17$, modifying the last $5$ blocks. For FLUX.1-dev, we edit from timestep $2$ to $11$ across the last $6$ blocks. For video editing with CogVideoX-2B, the official implementation is used. During editing, key ($\bm{K}$)and value ($\bm{V}$) tokens from the source branch are copied into the attention layers of the target branch.
  • FlowEdit (Kulikov et al., 2024): We adopt the official image editing setting for FLUX-1-dev. For SD3, we rescale the n_maxn\_max factor to account for the change in inference steps. Specifically, since the number of steps has changed from 50 to 28, we set $n\_max=28\div 50\times 33\approx 18$, while keeping all other parameters unchanged.
  • UniEdit-Flow (Jiao et al., 2025): The official version supports SD3 and FLUX.1-dev but provides the $\omega$ parameter only for CFG = $1$. Following the similarity transformation introduced in the paper, we use $\omega=5\div 7.5\approx 0.6$ and set $\alpha=0.6$ for SD3, $\alpha=0.85$ for FLUX.1-dev, and $\alpha=0.8$ for video generation, which yields comparable performance to the original.
  • FireFlow (Deng et al., 2025): In SD3, we observe that it is difficult to select a suitable end timestep for FireFlow, as setting it too high often leads to artifacts and generation failures, as shown in Fig. 9. To mitigate quality degradation caused by excessive editing steps, we restrict editing to timesteps 0 through 3 across all blocks in SD3. In contrast, FLUX.1-dev does not exhibit this issue, and using the same range (timesteps 0 to 3) yields results comparable to those reported in the original paper. For video generation, the editing ends at timestep $9$. During editing, value tokens ($\bm{V}$) are copied from the source to the target.
  • RF-Solver (Wang et al., 2025): RF-Solver exhibits a similar issue to FireFlow when applied to SD3, where higher editing timesteps lead to artifacts and degraded generation quality, as illustrated in Fig. 9. To address this, we limit editing to timesteps 0 to 7 in the second half of the SD3 blocks. For FLUX.1-dev, we set the end timestep to 4, which yields stable results without noticeable artifacts. Value tokens ($\bm{V}$) are copied from the source to the target. For video generation, editing is performed up to timestep $9$.
  • SDEdit (Meng et al., 2022): We set $t_{0}=0.6$ and apply editing to both generated and real input images or videos. The same parameter is applied across SD3, FLUX.1-dev, and CogVideoX-2B.
  • FLUX.1 Kontext Max (Black Forest Labs, 2025): Results are obtained using the official API, with source images and instruction prompts taken from PIE-Bench (Ju et al., 2024).
  • GPT-4o Image Generation (OpenAI, 2025): Results are generated through the official client using the source images and instruction prompts from PIE-Bench.
A.4. ColorCtrl-Bench Construction
Here, we provide details on the construction of ColorCtrl-Bench. Following a similar approach to that in Ju et al. (2024), we use GPT to generate a dataset of tuples, each consisting of a source prompt, a target prompt, a subject token, and an instruction, consistent with the format used in PIE-Bench. These tuples represent an image before and after a color editing operation, along with an instruction describing the transformation and a subject token indicating the object to be modified. The exact template is shown in Fig. 17.
A.5. Evaluation Details
For editing region masks, we first detect the target object with Grounded SAM 2 (Ren et al., 2024), using the “blended_word” in the PIE-Bench and ColorCtrl-Bench as the detection keyword. The raw mask is then dilated to compensate for the uncertainty of the boundary, slightly enlarging the edited region. A marginally eroded mask is still adequate for BG preservation metrics, and a marginally dilated foreground does not bias CLIP similarity scores. Hence, this dilation neither undermines the consistency check nor suffers from the boundary instability of Grounded SAM 2. For video evaluation, we compute each metric frame-wise and report the average over all frames as the final score for the video.
Appendix B More Results and Analysis
B.1. More Results of Image Editing
Method SD3 FLUX.1-dev
Canny BG Preservation $\text{Clip Similarity}$$\uparrow$ Canny BG Preservation $\text{Clip Similarity}$$\uparrow$
$\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited $\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited
FLUX.1 Kontext Max (Black Forest Labs, 2025) 0.8016 30.40 0.9152 28.23 26.83 0.8032 27.18 0.8854 27.56 26.24
GPT-4o Image Generation (OpenAI, 2025) 0.6988 21.69 0.8030 28.24 26.66 0.7727 23.05 0.8371 27.65 26.24
Ours 0.8775 38.16 0.9896 28.07 26.69 0.9324 37.96 0.9901 26.53 25.26
Table 6. Quantitative image results compared with commercial models on ColorCtrl-Bench.
Figure 10. Examples of image editing results.
Tab. 5 and Tab. 6 present quantitative results of image editing tasks on ColorCtrl-Bench, comparing our method against both training-free baselines and commercial models. Visual examples are shown in Fig. 13. Consistent with our findings on PIE-Bench, ColorCtrl outperforms other methods in maintaining object consistency and achieves accurate color edits with proper illumination and reflection. The similar conclusion drawn from this larger-scale benchmark further highlights the robustness and effectiveness of our approach.
Fig. 15 presents additional image editing results on PIE-Bench, comparing our method with both training-free and commercial models. In the first column, only ColorCtrl successfully preserves the texture of the rocks, while other methods either fail to change the color or introduce texture distortions. In the third column, FLUX.1 Kontext Max and GPT-4o alter the pose and fabric wrinkles, whereas other training-free methods fail to perform the intended edit.
Fig. 10 shows additional image results with FLUX.1-dev.
B.2. More Results of Video Editing
Fig. 14 and Fig. 11 present video editing results on ColorCtrl-Bench, while Fig. 16 shows additional results on PIE-Bench. Our method consistently outperforms all baselines, achieving superior color editing accuracy and better structural consistency across frames.
Figure 11. Additional qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “gray sofa” $\to$ “brown sofa”. Right: “brown suitcase” $\to$ “black suitcase”. Each shows three frames.
B.3. Limitations
The generation quality and the precision of text-guided localization in our method are fundamentally limited by the capabilities of the underlying generative models. As shown in Fig. 12, a failure case with SD3 demonstrates that the model fails to correctly detect “red trees” and instead modifies all trees, including green ones that should remain unchanged. Similarly, in the example with FLUX.1-dev, the model misinterprets “lipstick” and edits the casing rather than the lipstick itself. As foundation models continue to improve, we expect the performance and applicability of our method to advance accordingly.
Figure 12. Examples of failure cases. Left: results generated with SD3. Right: results generated with FLUX.1-dev.
Furthermore, editing real images and videos remains challenging due to the limitations of current inversion and reconstruction techniques. Although our method performs reliably on data that lies within the training distribution of the generative model, handling real-world inputs requires first mapping them accurately into the latent space of the model. This mapping process is still difficult and highly sensitive to the quality of the inversion.
Figure 13. Qualitative image results compared with training-free methods and commercial models on ColorCtrl-Bench. The left four columns are generated using FLUX.1-dev, while the right three are generated using SD3. Best viewed with zoom-in.
Figure 14. Qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “yellow bus” $\to$ “green yellow”. Right: “green backpack” $\to$ “yellow backpack”. Each shows three frames.
Figure 15. Additional qualitative image results compared with training-free methods and commercial models on PIE-Bench. The left three columns are generated using FLUX.1-dev, while the right two are generated using SD3. Best viewed with zoom-in.
Figure 16. Additional qualitative video results compared with training-free models on PIE-Bench. Left: “white shirt” $\to$ “blue shirt”. Right: “white lamb” $\to$ “blue lamb”. Each shows three frames.
Figure 17. The prompt used to generate ColorCtrl-Bench with GPT.